OpenStudy (anonymous):

If f(x) = x-2/2, then what is f–1(x) ? 2x + 2 2x – 1 2x – 2 1/2x – 1

OpenStudy (anonymous):

First thing you do is replace f(x) with y...\[y = \frac{x-2}{2}\] Switch the x's and y's \[x = \frac{y-2}{2}\] Solve for y\[y = 2x + 2\] replace y with \[f^{-1}(x)\] So the final answer is \[f^{-1}(x) = 2x + 2\]